# Channel Knowledge Map-assisted Dual-domain Tracking and Predictive Beamforming for High-Mobility Wireless Networks

### v1

@@ V-B B-Domain Measurement Model
After receiving the reflected echo, the RSU can calculate the cross-ambiguity function (CAF) [32] between the received signal $\boldsymbol{R}$ and the transmitted signal $\tilde{\boldsymbol{S}}$ to estimate the delay and Doppler shift of multiple paths. Following (2), a peak occurs at the matched filter output when $\tau$ and $\mu$ are perfectly tuned to the corresponding delay and Doppler shifts of the $i$-th path, respectively, i.e., $\{\hat{l}_{i},\hat{k}_{i}\}={\text{arg max}}_{l_{i},k_{i}}||\boldsymbol{R}\boldsymbol{\Lambda}_{k_{i}}^{H}\boldsymbol{S}_{l_{i}}^{H}||_{F}^{2},i=1,\dots,P$.
@@
$\small\boldsymbol{R}_{i}=\underbrace{\sum_{j=1}^{P}\beta_{j}\boldsymbol{b}(\theta_{j})\boldsymbol{a}^{H}(\theta_{j})\boldsymbol{F}{\boldsymbol{S}_{l_{j}}}{\boldsymbol{\Lambda}_{k_{j}}}\boldsymbol{\Lambda}_{k_{i}}^{H}\boldsymbol{S}_{l_{i}}^{H}}_{\text{Signal and interference: }\boldsymbol{R}_{i}^{\prime}}+\underbrace{\boldsymbol{Z}\boldsymbol{\Lambda}_{k_{i}}^{H}\boldsymbol{S}_{l_{i}}^{H}}_{\text{Noise: }\boldsymbol{R}_{i}^{\prime\prime}},$ (16)
where $\boldsymbol{R}_{i}^{\prime}$ can be approximated as $\boldsymbol{R}_{i}^{\prime}\approx L\beta_{i}\boldsymbol{b}(\theta_{i})\boldsymbol{a}^{H}(\theta_{i})\boldsymbol{F}$ based on the assumptions above. ∎
@@
Lemma 1 enables the RSU to separate the received signal from different paths for AoA estimation using the estimated delay and Doppler shift. According to (15), the received signal associated with the $i$-th path is rewritten as: $\boldsymbol{r}_{i}\triangleq{\text{vec}}(\boldsymbol{R}_{i})=L\beta_{i}(\boldsymbol{F}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i})+(\boldsymbol{S}_{l_{i}}^{*}\otimes\boldsymbol{Z}){\text{vec}}(\boldsymbol{\Lambda}_{k_{i}}^{H})\in\mathbb{C}^{N_{s}N_{r}\times 1}$. Given the AoA of the $i$-th path $\theta_{i}$, we can prove that $\boldsymbol{r}_{i}\sim\mathcal{CN}(L\beta_{i}(\boldsymbol{F}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i}),\sigma_{z}^{2}L\boldsymbol{I}_{N_{s}N_{r}})$ and the likelihood function of $\theta_{i}$ is given by
$\small p\left(\boldsymbol{r}_{i}|\theta_{i}\right)=\frac{1}{(L\pi\sigma_{z}^{2})^{N_{s}N_{r}}}\exp\{-\frac{1}{2\sigma_{z}^{2}}\bar{\boldsymbol{r}}_{i}^{H}\bar{\boldsymbol{r}}_{i}\}$ (17)
@@ V-C B-Domain Tracking
$\small\hat{\theta}_{i}[\eta+1|\eta]=\arg\max\operatorname{Pr}(\theta_{i}[\eta+1|\eta]).$ (18)
Based on the B-Domain soft state prediction $\boldsymbol{p}({\theta}_{i}[\eta|\eta-1])$ in the $[\eta-1]$-th time slot and new measurement $\boldsymbol{R}_{i}[\eta]$ in (15), we can update the AoA of the $i$-th path in $[\eta]$-th time slot employing the maximum a posteriori (MAP) principle, i.e.,
@@ VI CKM-assisted Predictive BF
In this section, we first derive the CRB of multipath AoA estimation and then design predictive BF for the subsequent time slot to minimize the maximum CRB for AoA estimation of each path, exploiting the channel a priori information obtained from the previous time slot. It is worth noting that accurate beam tracking benefits both target tracking and communication, which motivates us to design predictive BF for optimizing the beam tracking performance in the next time slot. The trade-off between dual-domain tracking and communication performance is a critical issue of the considered system. However, C-Domain tracking performance depends on B-Domain tracking performance and thus could be a complicated function of predictive BF strategy. Moreover, optimizing communication performance needs to consider the uncertainty of predicted CSI based on the predicted AoA for each path, which would lead to a complicated optimization problem. As a first attempt to propose a dual-domain tracking framework, this paper only considers the beam tracking performance in predictive BF design and the optimization of the trade-off between sensing and communication in the context of dual-domain tracking is left for our future work.
@@ VI-A CRB for Multi-path AoA Estimation
$\small\mathbb{E}[(\hat{\theta}_{i}-\theta_{i})^{2}]\geq{1}/{\mathcal{J}(\theta_{i})}\triangleq\operatorname{CRB}\left(\theta_{i}\right).$ (20)
Remark 1: The posterior Cramér-Rao bound (PCRB), which incorporates both the measured data and the a priori information of $\theta_{i}$, is commonly used to characterize the estimation performance in beam tracking frameworks [18]. However, as discussed in Fig. 4, a priori information might be unavailable for the third case of the condition of beam transition, i.e., non-stationary and unpredictable transition. Moreover, even for the first two types of AoA variation, the prior information obtained from TPM is numerical and thus cannot be integrated into PCRB derivation. Consequently, we adopt CRB as multipath AoA estimation performance metric. Note that in the $[\eta]$-th time slot, CRB in (20) depends on ground truth value of $\theta_{i}$ in the $[\eta+1]$-th time slot, which is unavailable. However, the B-Domain hard state prediction in (18) provides an AoA prediction $\hat{\theta}_{i}[\eta+1|\eta]$ which can be substituted into (20) to compute CRB. We note that this is one way to embed a prior information of $\theta_{i}$ into its predictive estimation error, as commonly adopted in the literature [18].
@@ VI-B Predictive BF Design
In $[\eta]$-th time slot, the predictive BF design to minimize the maximum predictive AoA estimation error $\mathrm{CRB}(\theta_{i}[\eta+1|\eta])$ is formulated as the following optimization problem
@@
Note that problem (22) is difficult to be solved globally due to the non-convex objective function. As a compromise, we propose a two-step optimization scheme with beam selection and multi-beam power allocation. In particular, the predictive BF matrix $\boldsymbol{F}[\eta+1|\eta]$ is defined as the product of an analog beam steering matrix $\boldsymbol{A}[\eta+1|\eta]\in\mathbb{C}^{N_{t}\times N_{s}}$ and a diagonal power allocation matrix $\boldsymbol{\Gamma}[\eta+1|\eta]\in\mathbb{R}^{N_{s}\times N_{s}}$, i.e., $\boldsymbol{F}[\eta+1|\eta]=\boldsymbol{A}[\eta+1|\eta]\boldsymbol{\Gamma}[\eta+1|\eta]$, where $\boldsymbol{\Gamma}[\eta+1|\eta]\triangleq\operatorname{diag}(\sqrt{\gamma_{1}[\eta+1|\eta]},\cdots,\sqrt{\gamma_{N_{s}}[\eta+1|\eta]})$, and $\gamma_{n_{s}}[\eta+1|\eta]$ denotes the power allocated to the $n_{s}$-th beam in the $[\eta+1]$-th slot.
@@ VI-B1 Beam Selection
where $\hat{\theta}_{n_{s}}[\eta+1|\eta]$ is the predicted AoA for the $n_{s}$-th path at the $[\eta]$-th time slot in (18) [^7] If $N_{s}<P$, the predicted AoAs with higher channel path gains should be selected for BF. Conversely, if $N_{s}\geq P$, the predicted AoAs for all $P$ paths can be selected. .
@@ VI-B2 Power Allocation
$\quad\cdot\left\{\left[\left(\boldsymbol{\Gamma}^{T}[\eta+1|\eta]\boldsymbol{A}^{T}[\eta+1|\eta]\right)\otimes\left(\operatorname{diag}(0,\ldots,N_{r}-1)\boldsymbol{b}(\theta_{i}[\eta+1|\eta])\right)\right]\boldsymbol{a}^{*}(\theta_{i}[\eta+1|\eta])\right\}.$
@@
where $c_{i,{n_{s}}}=\sum_{g=1}^{N_{r}}|\{\boldsymbol{c}(\theta_{i})\}_{({n_{s}}-1)N_{r}+g}|^{2}(g-1)^{2}$. After the proposed series of transformations, it can be verified that the power allocation problem (25) is convex and can be solved using off-the-shelf numerical convex programming tools, such as CVX [39].
@@ VII-A Simulation Parameters And Baseline Schemes
The RSU is positioned at the origin and the vehicle starts from the coordinates $[-20\mathrm{\leavevmode\nobreak\ m},10\mathrm{\leavevmode\nobreak\ m}]^{T}$ and moves along the road with an initial speed of $10\mathrm{\leavevmode\nobreak\ m/s}$. The time of interest is $T_{\max}=4\mathrm{\leavevmode\nobreak\ s}$ with a slot duration of $\Delta T=0.02\mathrm{\leavevmode\nobreak\ s}$. The carrier frequency $f_{c}$ is set to $30\mathrm{\leavevmode\nobreak\ GHz}$ and the sample interval $T_{p}$ is $10^{-8}$ seconds. Both the maximum delay and Doppler shift indices are set as $100$. The total power transmit $P_{\text{t}}$ is $16\mathrm{\leavevmode\nobreak\ W}$ and the angular grid size is set as $N_{\theta}=7200$. For the C-Domain state evolution noises, we set the standard deviations for position $\sigma_{q_{x}},\sigma_{q_{y}}$ are $10^{-3}\mathrm{\leavevmode\nobreak\ m}$, while $\sigma_{v}$ is $10^{-3}\mathrm{\leavevmode\nobreak\ m/s}$. The standard deviations for the delay and Doppler shift measurement noise are $\sigma_{\tau}=10^{-8}\mathrm{\leavevmode\nobreak\ s}$ and $\sigma_{\mu}=20\mathrm{\leavevmode\nobreak\ Hz}$, respectively. The standard deviation of angle measurement noise is $\sigma_{\cos(\theta)}=0.01$. For the stationary transition case, the TPM constant is $c_{\Pi}=0.3$ for simplicity, and AoA variation model parameters are $\xi=0.8$, $\epsilon=\hat{v}\sin(\hat{\theta}_{1})N_{\theta}/20$, and $\sigma_{\text{CKM}}=10^{-3}$.
@@
The ray tracing toolbox in MATLAB is leveraged to generate the ground-truth channel information [40]. A CKM is designed to record the $P=2$ strongest channel paths, denoted as path 1 and path 2, respectively. The CKM is constructed using the inverse distance weighting (IDW) of the $K$ nearest neighbors (KNN) approach with a limited set of 3200 ray tracing data samples [41]. The static shadowing caused by static scatterers in the propagation environment is assumed to emerge between the $140$-th and $175$-th time slots, whereas the random shadowing caused by a random obstacle appears with a certain blockage probability. Note that NLoS path is more likely to be blocked than LoS path in practice [42]. Therefore, we can define the blockage probability of the LoS path as $p_{\text{blk}}$ and the block probability of an NloS path as $1-(1-p_{\text{blk}})^{2}$ since an NLoS path can be treated as the concatenation of two LoS paths.
@@
In this section, a baseline method described in [4] is considered to demonstrate the benefits of the proposed framework. In particular, the baseline algorithm in [4] integrates the result of LoS path blockage detection with an EKF-based tracking framework where the position, speed, and LoS path AoA are jointly tracked. Note that the baseline is designed for solely based on the LoS channel model and fails to account for the information of NLoS paths. Therefore, we assume that the baseline scheme can perfectly separate the received signal of LoS path in the LoS-present case, which is utilized for tracking. If the LoS path is absent, the sensing receiver ignores the observation of the NLoS path and performs state prediction only to predict the position, speed, and LoS path AoA of the vehicle as considered in [4].

### v2

@@ V-B B-Domain Measurement Model
[…]
@@
[…]
To facilitate the derivation, without loss of generality, we assume $l_{j}\leq l_{i}$, $\forall j<i$. $\boldsymbol{R}_{i}^{\prime}$ can be simplified as:
$\boldsymbol{R}_{i}^{\prime}=$ $\beta_{i}\boldsymbol{b}(\theta_{i})\boldsymbol{a}^{H}(\theta_{i})\boldsymbol{F}\boldsymbol{S}\boldsymbol{S}^{H}$ (17)
$+\sum_{j\neq i}^{P}\beta_{j}\boldsymbol{b}(\theta_{j})\boldsymbol{a}^{H}(\theta_{j})\boldsymbol{F}{\boldsymbol{S}_{l_{j}}}\operatorname{diag}(\boldsymbol{0}_{l_{i}},\boldsymbol{\Lambda}_{ji},\boldsymbol{0}_{[\max(l_{j})-l_{j}]})\boldsymbol{S}_{l_{i}}^{H},$
where
$\boldsymbol{\Lambda}_{ji}$ $\triangleq\{\boldsymbol{\Lambda}^{\text{shift}}_{k_{i}}\}_{(l_{i}-l_{j}+1):L,(l_{i}-l_{j}+1):L}\{\boldsymbol{\Lambda}^{\text{shift}}_{k_{i}}\}_{1:(L-l_{i}+l_{j}),1:(L-l_{i}+l_{j})}^{H}.$
It follows that
$\boldsymbol{R}_{i}^{\prime}=$ $\beta_{i}\boldsymbol{b}(\theta_{i})\boldsymbol{a}^{H}(\theta_{i})\boldsymbol{F}\boldsymbol{S}\boldsymbol{S}^{H}$ (18)
$+\sum_{j\neq i}^{P}\beta_{j}\boldsymbol{b}(\theta_{j})\boldsymbol{a}^{H}(\theta_{j})\boldsymbol{F}\left(\sum_{l=l_{i}+1}^{L+l_{j}}\boldsymbol{x}_{L-l_{j}}\boldsymbol{x}_{L-l_{i}}^{H}\left\{\boldsymbol{\Lambda}_{ji}\right\}_{l,l}\right)$
$\approx$ $L\beta_{i}\boldsymbol{b}(\theta_{i})\boldsymbol{a}^{H}(\theta_{i})\boldsymbol{F}$
Therefore, for any $i=1,\cdots,P$, $\boldsymbol{R}_{i}$ is obtained by $\boldsymbol{R}_{i}\approx L\beta_{i}\boldsymbol{b}(\theta_{i})\boldsymbol{a}^{H}(\theta_{i})\boldsymbol{F}+\boldsymbol{Z}\boldsymbol{\Lambda}_{k_{i}}^{H}\boldsymbol{S}_{l_{i}}^{H}\in\mathbb{C}^{N_{r}\times N_{s}}$. ∎
@@
Lemma 1 enables the RSU to separate the received signal from different paths for AoA estimation using the estimated delay and Doppler shift. According to (15), the received signal associated with the $i$-th path is rewritten as: $\boldsymbol{r}_{i}\triangleq{\text{vec}}(\boldsymbol{R}_{i})=L\beta_{i}(\boldsymbol{F}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i})+(\boldsymbol{S}_{l_{i}}^{*}\otimes\boldsymbol{Z}){\text{vec}}(\boldsymbol{\Lambda}_{k_{i}}^{H})\in\mathbb{C}^{N_{s}N_{r}\times 1}$.
Lemma 2
Given the AoA of the $i$-th path $\theta_{i}$, $\boldsymbol{r}_{i}$ is a complex multivariate Gaussian random variable with $\boldsymbol{r}_{i}\sim\mathcal{CN}(L\beta_{i}(\boldsymbol{F}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i}),\sigma_{z}^{2}L\boldsymbol{I}_{N_{s}N_{r}})$.
Proof:
It is clear that the mean of $\boldsymbol{r}_{i}$ is $\mathbb{E}\left\{\boldsymbol{r}_{i}\right\}=L\beta_{i}(\boldsymbol{F}^{T}\otimes\boldsymbol{b}(\theta_{i}))\boldsymbol{a}^{*}(\theta_{i})$. Then, we proceed with the derivation of the covariance matrix for $\boldsymbol{r}_{i}$. Consider a $N_{r}$-dimensional vector denoted as $\tilde{\boldsymbol{z}}_{n}$, defined as $\tilde{\boldsymbol{z}}_{n}=\boldsymbol{Z}_{:,(l_{i}+1):(l_{i}+L)}\boldsymbol{D}^{H}_{n,:}$, where $\boldsymbol{Z}_{:,(l_{i}+1):(l_{i}+L)}\in\mathbb{C}^{N_{r}\times L}$ and $\boldsymbol{D}=\boldsymbol{S}\boldsymbol{\Lambda}^{\text{shift}}_{k_{i}}\in\mathbb{C}^{N_{s}\times L}$.
Consequently, we have
$\left\{\tilde{\boldsymbol{z}}_{n}\right\}_{m}$ $=\boldsymbol{Z}_{m,(l_{i}+1):(l_{i}+L)}\boldsymbol{D}^{H}_{n,:}$ (19)
$=\sum_{l=1}^{L}\boldsymbol{Z}_{m,(l_{i}+l)}\boldsymbol{S}^{*}_{n,l}e^{-j2\pi\mu_{i}lT_{p}},$
and its variance is given by $\sigma^{2}\left(\left\{\tilde{\boldsymbol{z}}_{n}\right\}_{m}\right)=\sigma^{2}_{z}\sum_{l=1}^{L}|\boldsymbol{S}^{*}_{n,l}e^{-j2\pi\mu_{i}lT_{p}}|^{2}=\sigma^{2}_{z}L$. Given that $\boldsymbol{Z}_{m,(l_{i}+l)}$ and $\boldsymbol{Z}_{m^{\prime},(l_{i}+l^{\prime})}$ are independent and identically distributed, i.e., $\mathbb{E}\{\boldsymbol{Z}_{m,(l_{i}+l)}\boldsymbol{Z}^{*}_{m^{\prime},(l_{i}+l^{\prime})}\}=\delta(m-m^{\prime})$. Therefore, it is sufficient to demonstrate that the elements $\left\{\tilde{\boldsymbol{z}}_{n}\right\}_{m}$ are independent and identically distributed (i.i.d.):
$\mathbb{E}\left\{\left\{\tilde{\boldsymbol{z}}_{n}\right\}_{m}\left\{\tilde{\boldsymbol{z}}_{n^{\prime}}\right\}^{*}_{m^{\prime}}\right\}=$ $\sum_{l=1}^{L}\sum_{l^{\prime}=1}^{L}\mathbb{E}\left\{\boldsymbol{Z}_{m,(l_{i}+l)}\boldsymbol{Z}^{*}_{m^{\prime},(l_{i}+l^{\prime})}\right\}$ (20)
$\cdot\boldsymbol{S}^{*}_{n,l}e^{-j2\pi\mu_{i}lT_{p}}\boldsymbol{S}_{n^{\prime},l^{\prime}}e^{j2\pi\mu_{i}l^{\prime}T_{p}}$
$=$ $\sum_{l=1}^{L}\sum_{l^{\prime}=1}^{L}\delta(m-m^{\prime})\delta(l-l^{\prime})$
$=$ $\delta(m-m^{\prime})\delta(n-n^{\prime})$
Based on Lemma 2, the likelihood function of $\theta_{i}$ is obtained as
$p\left(\boldsymbol{r}_{i}|\theta_{i}\right)=\frac{1}{(L\pi\sigma_{z}^{2})^{N_{s}N_{r}}}\exp\{-\frac{1}{2\sigma_{z}^{2}}\bar{\boldsymbol{r}}_{i}^{H}\bar{\boldsymbol{r}}_{i}\}$ (21)
@@ V-C B-Domain Tracking
$\hat{\theta}_{i}[\eta+1|\eta]=\arg\max\operatorname{Pr}(\theta_{i}[\eta+1|\eta]).$ (22)
Based on the B-Domain soft state prediction $\boldsymbol{p}({\theta}_{i}[\eta|\eta-1])$ in the $\eta-1]-$th time slot and new measurement $\boldsymbol{R}_{i}[\eta]$ in (15), we can update the AoA of the $i$-th path in $\eta]-$th time slot employing the maximum a posteriori (MAP) principle, i.e.,
@@ VI CKM-assisted Predictive BF
In this section, we first derive the CRB of multipath AoA estimation and then design predictive BF for the subsequent time slot to minimize the maximum CRB for AoA estimation of each path, exploiting the channel a priori information obtained from the previous time slot. It is worth noting that accurate B-Domain tracking benefits both target tracking and communication, which motivates us to design predictive BF for optimizing the B-Domain tracking performance in the next time slot. The trade-off between dual-domain tracking and communication performance is a critical issue of the considered system. However, C-Domain tracking performance depends on B-Domain tracking performance and thus could be a complicated function of predictive BF strategy. Moreover, optimizing communication performance needs to consider the uncertainty of predicted CSI based on the predicted AoA for each path, which would lead to a complicated optimization problem. As a first attempt to propose a dual-domain tracking framework, this paper only considers the B-Domain tracking performance in predictive BF design and the optimization of the trade-off between sensing and communication in the context of dual-domain tracking is left for our future work.
@@ VI-A CRB for Multi-path AoA Estimation
$\mathbb{E}[(\hat{\theta}_{i}-\theta_{i})^{2}]\geq{1}/{\mathcal{J}(\theta_{i})}\triangleq\operatorname{CRB}\left(\theta_{i}\right).$ (24)
Remark 2: The posterior Cramér-Rao bound (PCRB), which incorporates both the measured data and the a priori information of $\theta_{i}$, is commonly used to characterize the estimation performance in beam tracking frameworks [24]. However, as discussed in Fig. 4, a priori information might be unavailable for the third case of the condition of beam transition, i.e., non-stationary and unpredictable transition. Moreover, even for the first two types of AoA variation, the prior information obtained from TPM is numerical and thus cannot be integrated into PCRB derivation. Consequently, we adopt CRB as multipath AoA estimation performance metric. Note that in the $\eta]-$th time slot, CRB in (24) depends on ground truth value of $\theta_{i}$ in the $\eta+1]-$th time slot, which is unavailable. However, the B-Domain hard state prediction in (22) provides an AoA prediction $\hat{\theta}_{i}[\eta+1|\eta]$ which can be substituted into (24) to compute CRB. We note that this is one way to embed a prior information of $\theta_{i}$ into its predictive estimation error, as commonly adopted in the literature [24].
@@ VI-B Predictive BF Design
In $\eta]-$th time slot, the predictive BF design to minimize the maximum predictive AoA estimation error $\mathrm{CRB}(\theta_{i}[\eta+1|\eta])$ is formulated as the following optimization problem
@@
Note that problem (26) is difficult to be solved globally due to the non-convex objective function. As a compromise, we propose a two-step optimization scheme with beam selection and multi-beam power allocation. In particular, the predictive BF matrix $\boldsymbol{F}[\eta+1|\eta]$ is defined as the product of an analog beam steering matrix $\boldsymbol{A}[\eta+1|\eta]\in\mathbb{C}^{N_{t}\times N_{s}}$ and a diagonal power allocation matrix $\boldsymbol{\Gamma}[\eta+1|\eta]\in\mathbb{R}^{N_{s}\times N_{s}}$, i.e., $\boldsymbol{F}[\eta+1|\eta]=\boldsymbol{A}[\eta+1|\eta]\boldsymbol{\Gamma}[\eta+1|\eta]$, where $\boldsymbol{\Gamma}[\eta+1|\eta]\triangleq\operatorname{diag}(\sqrt{\gamma_{1}[\eta+1|\eta]},\cdots,\sqrt{\gamma_{N_{s}}[\eta+1|\eta]})$, and $\gamma_{n_{s}}[\eta+1|\eta]$ denotes the power allocated to the $n_{s}$-th beam in the $\eta+1]-$th slot.
@@ VI-B1 Beam Selection
where $\hat{\theta}_{n_{s}}[\eta+1|\eta]$ is the predicted AoA for the $n_{s}$-th path at the $\eta]-$th time slot in (22). If $N_{s}<P$, the predicted AoAs with higher channel path gains should be selected for BF. Conversely, if $N_{s}\geq P$, the predicted AoAs for all $P$ paths can be selected. This paper proposes to select the strongest paths to improve the dual-domain tracking performance. For B-Domain tracking, selecting the strongest paths typically results in higher received SNR, leading to more reliable AoA estimates and prediction, which are crucial for maintaining beam alignment. For C-Domain tracking, the Fisher information for target state estimation is dominated by the strongest paths. Specifically, the LoS path enables an analytical measurement model derived from its deterministic geometric relationship associated with the target state, providing high-quality measurements, while the strongest NLoS paths become critical in the LoS-absent case, enabling continuous tracking with the assistance of the CKM, which serves as a numerical measurement model.
@@ VI-B2 Power Allocation
$\qquad\cdot\left\{\left[\left(\boldsymbol{\Gamma}^{T}[\eta+1|\eta]\boldsymbol{A}^{T}[\eta+1|\eta]\right)\otimes\left(\operatorname{diag}(0,\ldots,N_{r}-1)\boldsymbol{b}(\theta_{i}[\eta+1|\eta])\right)\right]\boldsymbol{a}^{*}(\theta_{i}[\eta+1|\eta])\right\}.$
@@
where $c_{i,{n_{s}}}=\sum_{g=1}^{N_{r}}|\{\boldsymbol{c}(\theta_{i})\}_{({n_{s}}-1)N_{r}+g}|^{2}(g-1)^{2}$. After the proposed series of transformations, it can be verified that the power allocation problem (29) is convex and can be solved using off-the-shelf numerical convex programming tools, such as CVX [11]. Problem (29) is a linear programming problem with $N_{s}+1$ variables and $(P\times N_{s})+1$ linear constraints, which can be solved in polynomial time using interior-point methods with a complexity of $O((P+N_{s})^{3.5})$ [15]. For typical parameter settings in our simulations, this process requires fewer than $10^{4}$ floating-point operations, which can typically be executed within a few milliseconds, depending on the processor used, and is well below the slot duration $\Delta T$.
Remark 3: It is worth noting that the beam selection step defines both the optimization space and landscape for the subsequent power allocation design and thus directly determines the power distribution among the selected beams. In (29), the beam selection determines the number of optimization variables $\gamma_{1},\ldots,\gamma_{N_{s}}$ and hence the problem dimensionality. Power allocation then optimizes the power distribution among the selected beams to minimize the maximum AoA estimation error across paths. Furthermore, the coefficients $c_{i,{n_{s}}}$ of (29) depend jointly on the transmit and receive steering vectors of the RSU as well as the beam selection matrix $\boldsymbol{A}$. In turn, power allocation affects the updated and predicted AoA in the next time slot and then indirectly affects the beam selection.
Remark 4: The formulation of (25) can be generalized to a joint optimization problem by introducing a communication constraint, such as constraints on the minimum achievable rate or maximum outage probability, based on the predicted CSI which implies the sensing-communication tradeoff specifically. A more stringent communication constraint (e.g., a higher minimum rate requirement) tends to allocate more transmit power to communication, which reduces the power concentrated along the predicted AoAs and consequently increases the CRB for AoA estimation. Conversely, a less stringent communication constraint allows more power concentrated on the predicted AoAs, thereby minimizing AoA estimation errors. A joint optimization of predictive beamforming for both sensing and communication represents an important direction for our future work.
@@ VII-A Simulation Parameters And Baseline Schemes
The RSU is positioned at the origin and the vehicle starts from the coordinates $20\mathrm{\penalty 10000\ m},10\mathrm{\penalty 10000\ m}]^{T}$ and moves along the road with an initial speed of $10\mathrm{\penalty 10000\ m/s}$. The time of interest is $T_{\max}=4\mathrm{\penalty 10000\ s}$ with a slot duration of $\Delta T=0.02\mathrm{\penalty 10000\ s}$. The carrier frequency $f_{c}$ is set to $30\mathrm{\penalty 10000\ GHz}$ and the sample interval $T_{p}$ is $10^{-8}$ seconds. Both the maximum delay and Doppler shift indices are set as $100$. The total power transmit $P_{\text{t}}$ is $16\mathrm{\penalty 10000\ W}$ and the angular grid size is set as $N_{\theta}=7200$. For the C-Domain state evolution noises, we set the standard deviations for position $\sigma_{q_{x}},\sigma_{q_{y}}$ are $10^{-3}\mathrm{\penalty 10000\ m}$, while $\sigma_{v}$ is $10^{-3}\mathrm{\penalty 10000\ m/s}$. The standard deviations for the delay and Doppler shift measurement noise are $\sigma_{\tau}=10^{-8}\mathrm{\penalty 10000\ s}$ and $\sigma_{\mu}=20\mathrm{\penalty 10000\ Hz}$, respectively. The standard deviation of angle measurement noise is $\sigma_{\cos(\theta)}=0.01$ and AoA variation model parameters are $\xi=0.8$, $\epsilon=\hat{v}\sin(\hat{\theta}_{1})N_{\theta}/20$, and $\sigma_{\text{CKM}}=10^{-3}$. Each cumulative distribution function (CDF) curve is obtained from 100 Monte Carlo runs.
@@
The ray tracing toolbox in MATLAB is leveraged to generate the ground-truth channel information [33]. A CKM is designed to record the $P=2$ strongest channel paths, denoted as path 1 and path 2, respectively [^7] Note that we consider two paths since the simulated environment exhibits two dominant propagation paths that capture the major multipath structure of the considered scenario. Therefore, we set $P=2$ in our simulations. Moreover, the proposed scheme is applicable to the case with more propagation paths.. The CKM is constructed adopting the IDW of the $K$ nearest neighbors (KNN) approach with a limited set of 3200 ray tracing data samples, as commonly adopted in the literature [4, 6, 19]. The selected sample size achieves a spatial resolution of approximately $0.1$ m in the considered area, which is comparable to the vehicle displacement between consecutive slots ($v\Delta T=0.2$ m). The temporally correlated blockage caused by static and large size obstacles in the propagation environment is assumed to emerge between the $140$-th and $175$-th time slots, whereas random blockage events caused by transient obstacles appear with a certain blockage probability. Note that NLoS path is more likely to be blocked than LoS path in practice [8]. Therefore, we can define the blockage probability of the LoS path as $p_{\text{blk}}$ and the block probability of an NLoS path as $1-(1-p_{\text{blk}})^{2}$ since an NLoS path can be treated as the concatenation of two LoS paths. We note that LoS-absent happens when the LoS path is blocked. Besides, the blockage of NLoS path is not treated as a LoS-absent case.
Figure 5: Average AoA tracking errors with respect to $c_{\Pi_{i}}[\eta]$ for all paths and slots with $\sigma_{z}^{2}=10^{-6}\mathrm{\penalty 10000\ W}$, $N_{t}=N_{r}=32$, and $p_{\text{blk}}=0.15$.
For the stationary transition case, the TPM constant is set to $c_{\Pi_{i}}[\eta]=0.6$ for all paths and time slots. This value was determined based on the observed relationship between $c_{\Pi_{i}}$ and tracking accuracy observed in Fig. 5, which shows that the average AoA tracking error first decreases and then increases as $c_{\Pi_{i}}$ grows, particularly for path 1. The initial improvement in AoA tracking accuracy arises because a moderate $c_{\Pi_{i}}$ leverages the CKM prior in the angular TPM in (12) to enhance multipath AoA estimation. However, the CKM prior is inherently imperfect, since its input position is predicted rather than measured, which introduces a mismatch between the predicted and actual AoA. The C-Domain tracking error, together with the uncertainty of CKM construction, further reduces B-Domain tracking accuracy. Therefore, the average AoA tracking error increases with further increasing $c_{\Pi_{i}}$.
@@
In this section, a baseline method described in [53] is considered to demonstrate the benefits of the proposed framework. In particular, the baseline algorithm in [53] integrates the result of LoS path blockage detection with an EKF-based tracking framework where the position, speed, and LoS path AoA are jointly tracked. Note that the baseline is designed solely based on the LoS channel model and fails to account for the information of NLoS paths. Therefore, we assume that the baseline scheme can perfectly separate the received signal of LoS path in the LoS-present case, which is utilized for tracking. If the LoS path is absent, the sensing receiver ignores the observation of the NLoS path and performs state prediction only to predict the position, speed, and LoS path AoA of the vehicle as considered in [53].